OpenStudy (anonymous):

Stan's Subaru travels 280 miles on an average speed. If the car had gone 5 mph faster, the trip would have taken 1 hour less. Find the average speed. Help!

OpenStudy (perl):

d = r*t 280 = r * t 280 = (r + 5) * (t-1)

OpenStudy (anonymous):

that's it?

OpenStudy (perl):

now you can solve by substitution

OpenStudy (perl):

they both equal to 280

OpenStudy (perl):

r * t = (r + 5) * (t-1)

OpenStudy (anonymous):

so r*t=(180+5) (5-1)?

OpenStudy (anonymous):

I ment (280+5)*(5-1)